OpenStudy (anonymous):

how many real number solutions does the equation have? 1.-7x^2+6x+3=0 2.7x^2+8x+5=0

OpenStudy (anonymous):

6²-4(-7)(3) the discriminant of the 1st equation is positive, so therefore there will be 2 real number solutions. 8²-(4)(7)(5) = 64-140= not a real root zero real solutions. Do you need to find the solutions?

OpenStudy (anonymous):

no thanks

OpenStudy (anonymous):

Alright, you welcome, and good luck! Also, if you ever need... http://openstudy.com/users/student_basil#/updates/53542b8be4b066a1cac3b1a0